OpenStudy (anonymous):

Find the equation of the line through (4/3,-2) parallel to the y-axis.

OpenStudy (ganpat):

equation of line is y = mx + c where slope of the line m= (y2-y1)/(x2-x1) Question says, it is parallel to Y axis, i.e. it doesn't have any y intercept so two points for you will be (4/3,-2) and (4/3,0)

OpenStudy (ganpat):

find slope, substitute in given equation along with (x,y) which will be (4/3, -2) and calculate C.. Then, substitute value of m and c in equation y = mx + c.. that will be the answer..